# Analog Computing for Signal Processing and Communications – Part I: Computing with Microwave Networks

### v2

@@ Abstract
Index Terms:
Analog computing, Kalman filter, linear minimum mean square error (LMMSE) estimator, matrix inversion, microwave networks.
@@ I Introduction
Analog computers based on optical systems, or optical analog computers, leverage the speed and parallelism of light to process data, represented as optical signals, at an exceptionally high rate [4, 5]. Optical analog computers have been designed to efficiently perform linear operations [10] as well as compute the inverse of a given matrix [11]. More complicated tasks have also been performed through optical analog computers, specifically related to image processing and machine learning inference, as reviewed in [12]. In [13], a 2-dimensional spatial differentiator has been built to transform the input optical signal into its second-order derivative, enabling edge detection within an image. Machine learning inference in the analog domain has been demonstrated in [14], where the functioning of a neural network (NN) has been emulated via the propagation of an optical signal through multiple diffractive layers, each acting as a layer of a NN. Diffractive optics has also been used to build a hybrid optical-electronic convolutional NN, in which the convolution operation is performed in the analog domain through a phase mask [15]. Although optical analog computers have demonstrated significant advantages in low-power and ultra-fast processing, they typically rely on large-scale architectures, as their size cannot be reduced below the operational wavelength.
@@
To realize miniaturized analog computers at scales smaller than the operational wavelength, metamaterials have emerged as artificial materials designed to have unusual electromagnetic (EM) properties [6, 7]. Because of their artificially induced properties, metamaterials allow us to manipulate EM signals within sub-wavelength scales. Analog computers based on metamaterials have demonstrated the possibility of solving mathematical equations [16], including multiple equations in parallel exploiting input signals at different frequencies [17]. Particular interest has been dedicated to the 2-dimensional counterpart of metamaterials, namely metasurfaces, further reducing the dimensions of the computing devices. Metasurfaces can be engineered to apply specific transfer functions to the spatial and frequency components of the incident wave, including differentiation, integration, and convolution [6, 7]. To perform these mathematical operations, various metasurfaces technologies have been developed, such as space-time digital metasurfaces [18] and nonlocal metasurfaces having an angular-dependent response [19, 20]. Beyond linear transformations [21], metasurfaces have been designed to perform edge detection and pattern recognition, which are crucial operations in image processing [22, 23]. The capability of metasurfaces to perform matrix-vector multiplications in the analog domain has been exploited to implement programmable NNs performing recognition tasks [24] and reinforcement learning tasks [25]. Metasurfaces are also expected to play a crucial role in future wireless communications due to their ability to perform computations directly in the wave domain [26]. Possible applications of computing metasurfaces in wireless communications include performing sensing tasks and enhancing physical layer security [27, 28].
@@
Besides metamaterials and metasurfaces, resistive memory arrays have also emerged as a prominent technology for building modern analog computing devices [8, 9]. Particularly noteworthy are crosspoint memory arrays, in which resistive memory cells are located at the intersections of horizontal and vertical lines in a matrix shape. The values stored in these memory cells can naturally represent the entries of a matrix, enabling the memory array to perform matrix operations, such as matrix-vector multiplications, directly in the analog domain [29, 30]. This approach, referred to as in-memory computing or analog matrix computing, has demonstrated the capability to execute matrix-vector multiplications with arbitrarily high precision while consuming less energy compared to digital computing [31]. The parallelism offered by resistive memory arrays in computing matrix-vector multiplications has been harnessed to effectively implement and accelerate NNs [32, 33, 34]. Furthermore, analog matrix computing via resistive memory arrays has been used to advance future multiple-input multiple-output (MIMO) communications [35, 36].
@@
Several communities have developed modern analog computing solutions to accelerate signal processing operations [4]-[25], [29]-[34] as well as to foster wireless communications [26, 27, 28, 35, 36]. However, the efforts of these communities are often siloed, with limited inter-disciplinarity, and this fragmentation has led to a lack of a unified model and shared theoretical understanding. In this two-part paper, we aim to fill this gap, by focusing on analog computers that linearly process microwave signals. We develop the model of these analog computers by using rigorous microwave theory, characterize their input-output relationship, and explore their application to signal processing (in Part I) and wireless communications (in Part II). Specifically, the contributions of Part I of this paper are as follows.
@@
First, we derive a general model for an analog computer that processes microwave signals satisfying the superposition principle, which we denote as microwave linear analog computer (MiLAC). A MiLAC can be regarded as a microwave network having multiple ports, where input signals are applied and output signals are read. Once the input signals are applied, they propagate at light speed within the microwave network generating the output signals. In this way, specific operations can be instantly computed by the MiLAC, requiring no analog-digital conversion or digital computations. This is the first effort in providing a theoretical framework unifying previous works [4]-[36].
@@
Second, we characterize the output of a MiLAC as a function of the input signals and the impedance components constituting its microwave network. Specifically, we derive the expression of the output signals by using microwave theory, and regarding the MiLAC as a multiport microwave network. The obtained expressions involve matrix-matrix products and matrix inversions, which are computationally expensive operations in the digital domain, as their complexity scales cubically with the matrix size. Nevertheless, they can be instantly computed by a MiLAC once the microwave network is properly reconfigured and the input signals are applied. While it is known that specific matrix operations can be computed in the analog domain [4]-[36], we provide the general expression for the functions computable with a MiLAC.
@@
Third, we show that a MiLAC can be used as a computing device to efficiently calculate popular mathematical operations, namely the linear minimum mean square error (LMMSE) estimator and the inverse of a given matrix. Compared to a digital computer, a MiLAC can calculate these two operations with a reduction in computational complexity by a factor of $2.5\times 10^{4}$ and $5.5\times 10^{3}$, respectively. We define a set of primitive operations computable with a MiLAC, and show how they can be sequentially used jointly with digital operations to efficiently implement more sophisticated algorithms, such as the Kalman filter. Specifically, a Kalman filter can be implemented with a MiLAC by reducing the computational complexity by $1.1\times 10^{4}$ times. The sequential use of primitive operations computed in the analog domain is less explored in literature [4]-[36], and can further extend the set of possible functions implementable with analog computers.
@@
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\Re\{a\}$, $\Im\{a\}$, and $|a|$ refer to the real part, imaginary part, and absolute value of a complex scalar $a$, respectively. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $[\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $[\mathbf{A}]_{i,k}$, $[\mathbf{A}]_{:,k}$, and $\|\mathbf{A}\|_{F}$ refer to the transpose, conjugate transpose, $(i,k)$th element, $k$th column, and Frobenius norm of a matrix $\mathbf{A}$, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex number sets, respectively. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
Figure 1: Representation of a $P$-port MiLAC with input on $N$ ports, where $P=N+M$.
@@ II Model of a Microwave Linear Analog Computer
In this section, we introduce and model a class of analog computers, that we denote as MiLAC. We define a microwave analog computer as an analog computer that processes microwave EM signals, i.e., with frequencies from 300 MHz to 300 GHz [37]. In addition, we introduce a linear analog computer as an analog computer that satisfies the superposition principle, i.e., the output caused by multiple inputs is the sum of the outputs caused by each input individually. Thus, a MiLAC is defined as an analog computer that is both microwave and linear.
@@ II-A Microwave Network
The circuit implementing a MiLAC is in general a linear microwave network, following the definition of MiLAC. According to microwave network theory [37, Chapter 4], a linear microwave network having a certain number of ports $P$ can be characterized by its impedance matrix $\mathbf{Z}\in\mathbb{C}^{P\times P}$, which linearly relates the voltages and currents at the $P$ ports through
@@
where $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$ and $\mathbf{i}=[i_{1},\ldots,i_{P}]^{T}\in\mathbb{C}^{P\times 1}$ are the voltage and current vectors at the ports, with $v_{p}\in\mathbb{C}$ and $i_{p}\in\mathbb{C}$ being the voltage and current at the $p$th port, for $p=1,\ldots,P$. Alternatively, a linear microwave network with $P$ ports can also be represented by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$, given by $\mathbf{Y}=\mathbf{Z}^{-1}$, relating the currents and voltages at the $P$ ports as
@@
equivalently to (1) [37, Chapter 4]. Note that impedance and admittance matrices are equivalent representations of a linear microwave network, allowing its analysis regardless of its internal circuit topology.
@@ II-B Input Signals
Assume that the MiLAC receives the input on the first $N$ ports of its reconfigurable microwave network, with $N\leq P$. Note that this comes with no loss of generality, as the ports can always be reordered to ensure the input is applied on the first $N$ ports. In addition, if an input needs to be applied on an internal node of the microwave network, an additional port can be opened, making this model general. The input is applied by $N$ voltage sources with their series impedance $Z_{0}$, e.g., $Z_{0}=50~{}\Omega$, or, equivalently, series admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Denoting as $u_{n}\in\mathbb{C}$ the voltage of the $n$th voltage source, for $n=1,\ldots,N$, the input vector is given by $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Following Ohm’s law, the input voltage $u_{n}$ is related to the voltage $v_{n}$ and current $i_{n}$ at the $n$th port by $i_{n}=Y_{0}(u_{n}-v_{n})$, for $n=1,\ldots,N$, giving
@@ II-D Tunable Admittance Components
Figure 2: Representation of a $4$-port MiLAC with input on $2$ ports.
To realize a MiLAC with maximum flexibility, i.e., whose admittance matrix $\mathbf{Y}$ can be reconfigured to any complex matrix with dimension $P\times P$, we can implement its microwave network with $P^{2}$ tunable admittance components connecting each port to ground and to all other ports. Specifically, port $k$ is connected to ground through an admittance $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, and port $k$ is connected to port $i$ through an admittance $Y_{i,k}\in\mathbb{C}$, $\forall i\neq k$. An example of such a MiLAC with maximum flexibility is reported in Fig. 2, where we have $P=4$ ports and $N=2$ input signals.
This topology of tunable admittance components allows us to arbitrarily reconfigure the admittance matrix $\mathbf{Y}$, as it can be shown by following the definition of $\mathbf{Y}$ in (2). According to (2), $[\mathbf{Y}]_{i,k}$ is given by applying a voltage $v_{k}$ to port $k$, short-circuiting all the other ports, i.e., $v_{p}=0$, $\forall p\neq k$, measuring the current $i_{i}$ entering at port $i$, and computing the ratio
@@
for $i,k=1,\ldots,P$ [37, Chapter 4]. Following (5), the entries $\left[\mathbf{Y}\right]_{i,k}$ can be computed as a function of the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ by separately studying the two cases of off-diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i\neq k$ and diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i=k$. First, when $i\neq k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(a), since all the other admittance components in the microwave network do not impact the current $i_{i}$. Thus, we obtain the off-diagonal entry $\left[\mathbf{Y}\right]_{i,k}$ as $\left[\mathbf{Y}\right]_{i,k}=-Y_{i,k}$ following Ohm’s law, where the negative sign arises from the direction of the current $i_{i}$. Second, when $i=k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(b), since all the other admittance components do not influence $i_{k}$. In this case, the diagonal entry $\left[\mathbf{Y}\right]_{k,k}$ is given by the parallel of the $P$ admittance components $Y_{1,k},\ldots,Y_{P,k}$, i.e., $\left[\mathbf{Y}\right]_{k,k}=\sum_{p=1}^{P}Y_{p,k}$. By summarizing the two examined cases, we can write
@@
for $i,k=1,\ldots,P$, showing that the $P^{2}$ tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ can be adjusted to allow $\mathbf{Y}$ to assume any arbitrary value [^1] Equation (7) clarifies the direct relationship between the entries of the admittance matrix $\mathbf{Y}$ and the tunable components of the microwave network. Note that closed-form expressions are not available to relate the entries of the impedance matrix $\mathbf{Z}$ and the tunable components..
Note that it is also possible to implement a MiLAC with a reduced number of tunable admittance components. However, such a MiLAC will experience reduced flexibility since the microwave network cannot be reconfigured to take any arbitrary admittance matrix $\mathbf{Y}$, as a consequence of the reduced circuit complexity. In this work, we consider a MiLAC having maximum flexibility and leave the analysis of the trade-off between circuit complexity and flexibility as future research.
Figure 3: Circuit to be analyzed to compute (a) the off-diagonal entry $[\mathbf{Y}]_{i,k}$ and (b) the diagonal entry $[\mathbf{Y}]_{k,k}$ as a function of the tunable admittance components.

### v3

@@ I Introduction
Analog computers based on optical systems, or optical analog computers, leverage the intrinsic parallelism of light propagation to process data, represented as optical signals, at an exceptionally high rate [4, 5]. Optical analog computers have been designed to efficiently perform linear operations [10] as well as compute the inverse of a given matrix [11]. More complicated tasks have also been performed through optical analog computers, specifically related to image processing and machine learning inference, as reviewed in [12]. In [13], a 2-dimensional spatial differentiator has been built to transform the input optical signal into its second-order derivative, enabling edge detection within an image. Machine learning inference in the analog domain has been demonstrated in [14], where an optical system based on multiple diffractive layers was proposed to emulate the forward propagation of a neural network (NN). However, this system is purely linear, and as pointed out in [15], it lacks the nonlinear activations that are pivotal in NN. Diffractive optics has also been used to build a hybrid optical-electronic convolutional NN, in which the convolution operation is performed in the analog domain through a phase mask [16]. Although optical analog computers have demonstrated significant advantages in low-power and ultra-fast processing, they typically rely on large-scale architectures, as their size cannot be reduced below the operational wavelength.
@@
To realize miniaturized analog computers at scales smaller than the operational wavelength, metamaterials have emerged as artificial materials designed to have unusual electromagnetic (EM) properties [6, 7]. Because of their artificially induced properties, metamaterials allow us to manipulate EM signals within sub-wavelength scales. Analog computers based on metamaterials have demonstrated the possibility of solving mathematical equations [17], including multiple equations in parallel exploiting input signals at different frequencies [18]. Particular interest has been dedicated to the 2-dimensional counterpart of metamaterials, namely metasurfaces, further reducing the dimensions of the computing devices. Metasurfaces can be engineered to apply specific transfer functions to the spatial and frequency components of the incident wave, including differentiation, integration, and convolution [6, 7, 19]. To perform these mathematical operations, various metasurfaces technologies have been developed, such as space-time digital metasurfaces [20] and nonlocal metasurfaces having an angular-dependent response [21, 22]. Beyond linear transformations [23], metasurfaces have been designed to perform edge detection and pattern recognition, which are crucial operations in image processing [24, 25]. The capability of metasurfaces to perform matrix-vector multiplications in the analog domain has been exploited to implement programmable NNs [26], performing recognition tasks [27] and reinforcement learning tasks [28]. Metasurfaces are also expected to play a crucial role in future wireless communications due to their ability to perform computations directly in the wave domain [29]. Possible applications of computing metasurfaces in wireless communications include performing sensing tasks and enhancing physical layer security [30, 31].
@@
Besides metamaterials and metasurfaces, resistive memory arrays have also emerged as a prominent technology for building modern analog computing devices [8, 9]. Particularly noteworthy are crosspoint memory arrays, in which resistive memory cells are located at the intersections of horizontal and vertical lines in a matrix shape [32]. The values stored in these memory cells can naturally represent the entries of a matrix, enabling the memory array to perform matrix operations, such as matrix-vector multiplications and solving linear algebra problems, directly in the analog domain [33, 34, 35, 36]. This approach, referred to as in-memory computing or analog matrix computing, has demonstrated the capability to execute matrix-vector multiplications with arbitrarily high precision while consuming less energy compared to digital computing [37]. The parallelism offered by resistive memory arrays in computing matrix-vector multiplications has been harnessed to effectively implement and accelerate NNs [38, 39, 40] and image processing [41]. Furthermore, analog matrix computing via resistive memory arrays has been used to advance future multiple-input multiple-output (MIMO) communications [42, 43].
@@
Several communities have developed modern analog computing solutions to accelerate signal processing operations [4]-[28], [33]-[40] as well as to foster wireless communications [29, 30, 31, 42, 43]. However, the efforts of these communities are often siloed, with limited interdisciplinarity, and this fragmentation has led to a lack of a unified model and shared theoretical understanding. In this two-part paper, we aim to fill this gap, by focusing on analog computers that linearly process microwave signals. We develop the model of these analog computers by using rigorous microwave theory, characterize their input-output relationship, and explore their application to signal processing (in Part I) and wireless communications (in Part II). Specifically, the contributions of Part I of this paper are as follows.
@@
First, we investigate the fundamental computational limits of a reconfigurable microwave network made of linear components. To this end, we derive a general model for an analog computer that processes microwave signals satisfying the superposition principle, which we denote as microwave linear analog computer (MiLAC). A MiLAC can be regarded as a microwave network having multiple ports, where input signals are applied and output signals are read. Once the input signals are applied, they propagate at light speed within the microwave network, generating the output signals. In this way, specific operations can be instantly computed by the MiLAC, requiring no analog-digital conversion or digital computations. This is the first effort in providing a theoretical framework unifying previous works [4]-[43].
@@
Second, we characterize the output of a MiLAC as a function of both the input signals and the impedance components constituting its microwave network. Specifically, we derive the expression of the output signals by using microwave theory, and regarding the MiLAC as a multiport microwave network. The obtained expressions involve matrix-matrix products and matrix inversions, which are computationally expensive operations in the digital domain, as their complexity scales cubically with the matrix size. Nevertheless, they can be instantly computed by a MiLAC once the microwave network is properly reconfigured and the input signals are applied. While it is known that specific matrix operations can be computed in the analog domain [4]-[43], we provide the general expression for the functions computable with microwave signals by a linear network.
@@
Third, we show that a MiLAC can be used as a computing device to efficiently calculate popular mathematical operations, namely the linear minimum mean square error (LMMSE) estimator and the inverse of a given matrix. Compared to a digital computer, a MiLAC can calculate these two operations with a reduction in computational complexity by a factor of $2.5\times 10^{4}$ and $5.5\times 10^{3}$, respectively, when the input size is $8192$. We define a set of primitive operations computable with a MiLAC, and show how they can be sequentially used jointly with digital operations to efficiently implement more sophisticated algorithms, such as the Kalman filter. Specifically, a Kalman filter can be implemented with a MiLAC by reducing the computational complexity by $1.1\times 10^{4}$ times, with input size being $8192$. The sequential use of primitive operations computed in the analog domain is less explored in literature [4]-[43], and can further extend the set of possible functions implementable with analog computers.
@@
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\Re\{a\}$, $\Im\{a\}$, and $|a|$ refer to the real part, imaginary part, and absolute value of a complex scalar $a$, respectively. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $\mathbf{A}]_{i,k}$, $\mathbf{A}]_{:,k}$, and $\|\mathbf{A}\|_{F}$ refer to the transpose, conjugate transpose, $(i,k)$th element, $k$th column, and Frobenius norm of a matrix $\mathbf{A}$, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex number sets, respectively. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
Figure 1: Representation of a $P$-port MiLAC with input on NN ports, where $P=N+M$.
@@ II Model of a Microwave Linear Analog Computer
In this section, we introduce and model a class of analog computers, that we denote as MiLAC. We define a microwave analog computer as an analog computer that processes microwave EM signals, i.e., with frequencies from 300 MHz to 300 GHz [44]. In addition, we introduce a linear analog computer as an analog computer that satisfies the superposition principle, i.e., the output caused by multiple inputs is the sum of the outputs caused by each input individually. Thus, a MiLAC is defined as an analog computer that is both microwave and linear. In the following discussion, it is important to specify the microwave regime as all results are derived by applying microwave theory [44].
@@ II-A Microwave Network
The circuit implementing a MiLAC is in general a linear microwave network, following the definition of MiLAC. According to microwave network theory [44, Chapter 4], a linear microwave network having a certain number of ports $P$ can be characterized by its impedance matrix $\mathbf{Z}\in\mathbb{C}^{P\times P}$, which linearly relates the voltages and currents at the $P$ ports through
@@
where $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$ and $\mathbf{i}=[i_{1},\ldots,i_{P}]^{T}\in\mathbb{C}^{P\times 1}$ are the voltage and current vectors at the ports, with $v_{p}\in\mathbb{C}$ and $i_{p}\in\mathbb{C}$ being the voltage and current at the $p$th port, for $p=1,\ldots,P$.  [^1] Throughout the paper, voltage and current signals are narrowband basspand signals represented through their complex-valued baseband equivalent. For example, for the voltage at the $p$th port, $v_{p}\in\mathbb{C}$ is the baseband equivalent, while the actual passband signal is $\Re\{v_{p}e^{j2\pi f_{c}t}\}$, where $f_{c}$ is the carrier frequency. In other words, the complex number $v_{p}=Ae^{j\varphi}$ represents the signal $A\cos(2\pi f_{c}t+\varphi)$, or equivalently, the complex number $v_{p}=I+jQ$ represents the signal $I\cos(2\pi f_{c}t)-Q\sin(2\pi f_{c}t)$. Alternatively, a linear microwave network with $P$ ports can also be represented by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$, given by $\mathbf{Y}=\mathbf{Z}^{-1}$, relating the currents and voltages at the $P$ ports as
@@
equivalently to (1) [44, Chapter 4]. Impedance and admittance matrices are equivalent representations of a linear microwave network, allowing its analysis regardless of its internal circuit topology.
@@ II-B Input Signals
Assume that the MiLAC receives the input on the first $N$ ports of its reconfigurable microwave network, with $N\leq P$. This comes with no loss of generality, as the ports can always be reordered to ensure the input is applied on the first $N$ ports. In addition, if an input needs to be applied on an internal node of the microwave network, an additional port can be opened, making this model general. The input is applied by $N$ voltage sources with their series impedance $Z_{0}$, e.g., $Z_{0}=50~\Omega$, or, equivalently, series admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Denoting as $u_{n}\in\mathbb{C}$ the voltage of the $n$th voltage source, for $n=1,\ldots,N$, the input vector is given by $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Following Ohm’s law, the input voltage $u_{n}$ is related to the voltage $v_{n}$ and current $i_{n}$ at the $n$th port by $i_{n}=Y_{0}(u_{n}-v_{n})$, for $n=1,\ldots,N$, giving
@@ II-D Tunable Admittance Components
Figure 2: Representation of a 44-port MiLAC with input on 22 ports.
To realize a MiLAC with maximum flexibility, i.e., whose admittance matrix $\mathbf{Y}$ can be reconfigured to any complex matrix with dimension $P\times P$, we can implement its microwave network with $P^{2}$ tunable admittance components connecting each port to ground and to all other ports. We characterize these components through the lumped-element model. Specifically, port $k$ is connected to ground through an admittance $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, and port $k$ is connected to port $i$ through an admittance $Y_{i,k}\in\mathbb{C}$, $\forall i\neq k$. An example of such a MiLAC with maximum flexibility is reported in Fig. 2, where we have $P=4$ ports and $N=2$ input signals.
Figure 3: Circuit to be analyzed to compute (a) the off-diagonal entry $\mathbf{Y}]_{i,k}$ and (b) the diagonal entry $\mathbf{Y}]_{k,k}$ as a function of the tunable admittance components.
This topology of tunable admittance components allows us to arbitrarily reconfigure the admittance matrix $\mathbf{Y}$, as it can be shown by following the definition of $\mathbf{Y}$ in (2). According to (2), $\mathbf{Y}]_{i,k}$ is given by applying a voltage $v_{k}$ to port $k$, short-circuiting all the other ports, i.e., $v_{p}=0$, $\forall p\neq k$, measuring the current $i_{i}$ entering at port $i$, and computing the ratio
@@
for $i,k=1,\ldots,P$ [44, Chapter 4]. Following (5), the entries $\left[\mathbf{Y}\right]_{i,k}$ can be computed as a function of the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ by separately studying the two cases of off-diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i\neq k$ and diagonal entries $\left[\mathbf{Y}\right]_{i,k}$ with $i=k$. First, when $i\neq k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(a), since all the other admittance components in the microwave network do not impact the current $i_{i}$. Thus, we obtain the off-diagonal entry $\left[\mathbf{Y}\right]_{i,k}$ as $\left[\mathbf{Y}\right]_{i,k}=-Y_{i,k}$ following Ohm’s law, where the negative sign arises from the direction of the current $i_{i}$. Second, when $i=k$, the ratio in (5) can be determined by studying the circuit in Fig. 3(b), since all the other admittance components do not influence $i_{k}$. In this case, the diagonal entry $\left[\mathbf{Y}\right]_{k,k}$ is given by the parallel of the $P$ admittance components $Y_{1,k},\ldots,Y_{P,k}$, i.e., $\left[\mathbf{Y}\right]_{k,k}=\sum_{p=1}^{P}Y_{p,k}$. By summarizing the two examined cases, we can write
@@
for $i,k=1,\ldots,P$, showing that the $P^{2}$ tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ can be adjusted to allow $\mathbf{Y}$ to assume any arbitrary value [^2] Equation (7) clarifies the direct relationship between the entries of the admittance matrix $\mathbf{Y}$ and the tunable components of the microwave network. Similar relationships in closed-form are not available to relate the entries of the impedance matrix $\mathbf{Z}$ and the tunable components..
It is also possible to implement a MiLAC with a reduced number of tunable admittance components. However, such a MiLAC will experience reduced flexibility since the microwave network cannot be reconfigured to take any arbitrary admittance matrix $\mathbf{Y}$, as a consequence of the reduced circuit complexity. In this work, we consider a MiLAC having maximum flexibility and leave the analysis of the trade-off between circuit complexity and flexibility as future research.